# MotionCraft: Physics-based Zero-Shot Video Generation

### v1

@@ Abstract
Generating videos with realistic and physically plausible motion is one of the main recent challenges in computer vision. While diffusion models are achieving compelling results in image generation, video diffusion models are limited by heavy training and huge models, resulting in videos that are still biased to the training dataset. In this work we propose MotionCraft, a new zero-shot video generator to craft physics-based and realistic videos. MotionCraft is able to warp the noise latent space of an image diffusion model, such as Stable Diffusion, by applying an optical flow derived from a physics simulation. We show that warping the noise latent space results in coherent application of the desired motion while allowing the model to generate missing elements consistent with the scene evolution, which would otherwise result in artefacts or missing content if the flow was applied in the pixel space. We compare our method with the state-of-the-art Text2Video-Zero reporting qualitative and quantitative improvements, demonstrating the effectiveness of our approach to generate videos with finely-prescribed complex motion dynamics. Project page: https://mezzelfo.github.io/MotionCraft/.
**footnotetext: indicates equal contribution.
@@ 3.2 Physics-based zero shot video generation
[…]
@@
Based on the analysis presented in the previous section, we propose a novel zero-shot video generation method, named MotionCraft, where an image (real or generated), serving as a starting frame $I^{0}$, is animated according to a physical simulation, by means of a (possibly time-varying) optical-flow generator $\mathcal{W}$ in the noise latent space. The outcome is a video made of NN frames $I^{0},\dots,I^{N-1}$ that follows the motion prescribed by the physical simulation and evolves the content of the first frame coherently. Inspired by the previous observation, this animation is obtained by warping the noisy latent representation of an image in the latent diffusion space. Regarding the physics simulation for the optical flow generation, we use different libraries to simulate different physics, as explained in the experimental section, such as fluid dynamics, rigid motion and multi-agent systems. It is also possible, albeit not shown in this paper, to use animation software to generate the required optical flows.
@@
Fig. 3 illustrates an overview of MotionCraft highlighting the autoregressive generation of the video. At each iteration $f\geq 1$, the frame $I^{f}$ is generated using only the information contained in the first frame $I^{0}$ and the previous frame $I^{f-1}$. Given this Markovian structure, MotionCraft is characterized by $\mathcal{O}\left(1\right)$ space complexity and $\mathcal{O}\left(N\right)$ time complexity with respect to the total number NN of frames to be generated. More in detail, first, the two RGB frames $I^{0},I^{f-1}$ are encoded into the latent space and they are independently inverted with the reversed DDIM sampling scheme up to a fixed diffusion timestep $\tau$, obtaining $z_{\tau}^{0}$, and $z_{\tau}^{f-1}$, respectively. Then, the optical flow warping operator $\mathcal{W}^{f-1\to f}$ prescribed by the physical simulation is applied to $z_{\tau}^{f-1}$, obtaining $\zeta_{\tau}^{f}$. Finally, the next RGB frame $I^{f}$ is generated by performing $\tau$ steps of reverse diffusion using the DDIM sampling scheme with a novel cross-frame attention mechanism and a novel spatial noise map $\eta^{f}$ weighting technique, explained below. Furthermore, we exploit the classifier-free guidance (CFG) technique for generation proposed in [14], with $\mathcal{P}$ and $\mathcal{P}_{\emptyset}$ being the positive and negative prompt, respectively, and $\gamma>1$ being the strength of the CFG. More details can be found in the Appendix A.
@@
Algorithm 1 reports the pseudocode of MotionCraft. Lines 262-6 include the DDIM inversion up to timestep $\tau$. Starting current frame $I^{f-1}$ that was previously generated, in line 22 we embed it with the VAE encoder $\mathcal{E}$, obtaining $z^{f-1}_{0}$. Then we apply DDIM inversion on $z^{f-1}_{0}$ for $\tau$ timesteps (line 363-6). This involves the UNet with the standard self-attention (note the repetition of the noisy latent $z^{f-1}_{t}$) and the positive prompt $\mathcal{P}$. As briefly reported in [21], we have also experienced that DDIM inversion is not compatible with CFG; hence, during the inversion, we do not use the negative prompt $\mathcal{P}_{\emptyset}$. The resulting estimated noise is used in line 55 for applying the DDIM inversion step (note that the $\eta=0$, so pure DDIM is performed). Upon completion of the DDIM inversion process, we obtain $z_{\tau}^{f-1}$, the noisy latent corresponding to the frame $I^{f-1}$.
@@
In line 77, the optical flow warping operator $\mathcal{W}^{f-1\to f}$ is applied to the noisy latent of the current frame $z_{\tau}^{f-1}$ to obtain a new noisy latent $\zeta_{\tau}^{f}$ that will generate the successive frame. Finally, in lines 8148-14, the frame is generated. During this generation phase we use CFG to increase the quality of the generated images, hence also the negative prompt $\mathcal{P}_{\emptyset}$ is used. To create new content while preserving the original image, we propose two direct generalization of two known techniques: the multiple cross-frame attention (MCFA) mechanism and a spatial noise map weighting (Spatial-$\eta$).
@@
The MCFA technique generalizes the Cross Frame Attention (CFA) [20], as it enables the to-be-generated frame to attend to an arbitrary number of frames. We choose to attend to the first frame and the previous frame (as shown in lines 9109-10 of Alg. 1 and Fig. 3) to ensure long-range and short-range temporal consistency, respectively. MCFA intervenes in all the self-attention blocks of the SD UNet, by replacing the keys and values, that are originally computed from projections of the generating frame features, with the ones computed from the attended frames.
@@
We also propose Spatial-$\eta$ (line 1212), that is a novel technique that enables to choose, on a pixel-by-pixel basis, whether to use DDIM or DDPM as a sampling scheme. This enables the usage of DDPM in regions of the images where novel content should be created (for example, when a new part of an object is entering the scene), while using DDIM in the other regions to ensure consistency and determinism where the already-present content is just moving. Note that this spatial map $\eta^{f}$ can be obtained in multiple ways from the physical simulation. For example, $\eta^{f}$ can be set to 11 in regions of the image where the flow is not well-defined (pointing outside of the image boundaries) or in regions where the optical flow field has discontinuities.
@@ 4.1 Experimental setting
In this section, we show examples of video generation based on different physics simulations: rigid body motion, fluid dynamics and multi-agent systems. Given an optical flow, we apply it on the SD latent space using MotionCraft. Then, we compare our method to Text2Video-Zero [20] that, to the best of our knowledge, is the only diffusion-based zero-shot method for video generation.
@@ Appendix A Background
$\text{Cross-Frame-Attn(Q,K,V)}=\text{Softmax}\left(\frac{Q^{f}\cdot K^{1}}{\sqrt{d_{k}}}\right)V^{1}$ (3)
@@ Appendix B Extendend Ablation Study
In this section we show the remaining ablations for the scene Earth, and additional ablations on two new scenes: Dragons and Satellite Scan. The ablations for cross frame attention mechanism can be found in Figures 10 and 11. The ablations of the Spatial-$\eta$ are shown in Figures 12 and 13. Moreover, we also show the contribution of the inversion mechanism in Figures 14, 15, and 16.
@@ Appendix C Obstacles, Different Physics and Additional Visual Results
Figure 17: Fluid simulation: pouring drink. Top: MotionCraft; Bottom: T2V0 [20].
Figure 18: Smoke simulation: Evaporating man. Top: MotionCraft; Bottom: T2V0.
@@ Appendix F Implementation Details and Licenses
We used the following hyperparameters throughout the work if not explicitely said otherwise. We set $\tau=400$, the number of inference steps (both for DDIM inversion and for inverse diffusion) is set to 200200 and the used model is runwayml/stable-diffusion-v1-5 (license CreativeML Open RAIL-M). All our experiments are done on a single NVIDIA A6000 (48GB); video generation runs in minutes (1-5min) on a single GPU. Our provided code is available under MIT license. The Earth image is a composite of six separate orbits taken on January 23, 2012 by the Suomi National Polar-orbiting Partnership satellite (Credit: NASA/NOAA).

### v2

@@ Abstract
Generating videos with realistic and physically plausible motion is one of the main recent challenges in computer vision. While diffusion models are achieving compelling results in image generation, video diffusion models are limited by heavy training and huge models, resulting in videos that are still biased to the training dataset. In this work we propose MotionCraft, a new zero-shot video generator to craft physics-based and realistic videos. MotionCraft is able to warp the noise latent space of an image diffusion model, such as Stable Diffusion, by applying an optical flow derived from a physics simulation. We show that warping the noise latent space results in coherent application of the desired motion while allowing the model to generate missing elements consistent with the scene evolution, which would otherwise result in artefacts or missing content if the flow was applied in the pixel space. We compare our method with the state-of-the-art Text2Video-Zero reporting qualitative and quantitative improvements, demonstrating the effectiveness of our approach to generate videos with finely-prescribed complex motion dynamics.
**footnotetext: indicates equal contribution.footnotetext: Project page: https://mezzelfo.github.io/MotionCraft/.
@@ 3.2 Physics-based zero shot video generation
             $\hat{\epsilon}\leftarrow\hat{\epsilon}_{\emptyset}+\gamma(\hat{\epsilon}_{\mathcal{P}}-\hat{\epsilon}_{\emptyset})$ ;
@@
Based on the analysis presented in the previous section, we propose a novel zero-shot video generation method, named MotionCraft, where an image (real or generated), serving as a starting frame $I^{0}$, is animated according to a physical simulation, by means of a (possibly time-varying) optical-flow generator $\mathcal{W}$ in the noise latent space. The outcome is a video made of $N$ frames $I^{0},\dots,I^{N-1}$ that follows the motion prescribed by the physical simulation and evolves the content of the first frame coherently. Inspired by the previous observation, this animation is obtained by warping the noisy latent representation of an image in the latent diffusion space. Regarding the physics simulation for the optical flow generation, we use different libraries to simulate different physics, as explained in the experimental section, such as fluid dynamics, rigid motion and multi-agent systems. It is also possible, albeit not shown in this paper, to use animation software to generate the required optical flows.
@@
Fig. 3 illustrates an overview of MotionCraft highlighting the autoregressive generation of the video. At each iteration $f\geq 1$, the frame $I^{f}$ is generated using only the information contained in the first frame $I^{0}$ and the previous frame $I^{f-1}$. Given this Markovian structure, MotionCraft is characterized by $\mathcal{O}\left(1\right)$ space complexity and $\mathcal{O}\left(N\right)$ time complexity with respect to the total number $N$ of frames to be generated. More in detail, first, the two RGB frames $I^{0},I^{f-1}$ are encoded into the latent space and they are independently inverted with the reversed DDIM sampling scheme up to a fixed diffusion timestep $\tau$, obtaining $z_{\tau}^{0}$, and $z_{\tau}^{f-1}$, respectively. Then, the optical flow warping operator $\mathcal{W}^{f-1\to f}$ prescribed by the physical simulation is applied to $z_{\tau}^{f-1}$, obtaining $\zeta_{\tau}^{f}$. Finally, the next RGB frame $I^{f}$ is generated by performing $\tau$ steps of reverse diffusion using the DDIM sampling scheme with a novel cross-frame attention mechanism and a novel spatial noise map $\eta^{f}$ weighting technique, explained below. Furthermore, we exploit the classifier-free guidance (CFG) technique for generation proposed in [14], with $\mathcal{P}$ and $\mathcal{P}_{\emptyset}$ being the positive and negative prompt, respectively, and $\gamma>1$ being the strength of the CFG. More details can be found in the Appendix A.
@@
Algorithm 1 reports the pseudocode of MotionCraft. Lines $2-6$ include the DDIM inversion up to timestep $\tau$. Starting current frame $I^{f-1}$ that was previously generated, in line $2$ we embed it with the VAE encoder $\mathcal{E}$, obtaining $z^{f-1}_{0}$. Then we apply DDIM inversion on $z^{f-1}_{0}$ for $\tau$ timesteps (line $3-6$). This involves the UNet with the standard self-attention (note the repetition of the noisy latent $z^{f-1}_{t}$) and the positive prompt $\mathcal{P}$. As briefly reported in [21], we have also experienced that DDIM inversion is not compatible with CFG; hence, during the inversion, we do not use the negative prompt $\mathcal{P}_{\emptyset}$. The resulting estimated noise is used in line $5$ for applying the DDIM inversion step (note that the $\eta=0$, so pure DDIM is performed). Upon completion of the DDIM inversion process, we obtain $z_{\tau}^{f-1}$, the noisy latent corresponding to the frame $I^{f-1}$.
@@
In line $7$, the optical flow warping operator $\mathcal{W}^{f-1\to f}$ is applied to the noisy latent of the current frame $z_{\tau}^{f-1}$ to obtain a new noisy latent $\zeta_{\tau}^{f}$ that will generate the successive frame. Finally, in lines $8-14$, the frame is generated. During this generation phase we use CFG to increase the quality of the generated images, hence also the negative prompt $\mathcal{P}_{\emptyset}$ is used. To create new content while preserving the original image, we propose two direct generalization of two known techniques: the multiple cross-frame attention (MCFA) mechanism and a spatial noise map weighting (Spatial-$\eta$).
@@
The MCFA technique generalizes the Cross Frame Attention (CFA) [20], as it enables the to-be-generated frame to attend to an arbitrary number of frames. We choose to attend to the first frame and the previous frame (as shown in lines $9-10$ of Alg. 1 and Fig. 3) to ensure long-range and short-range temporal consistency, respectively. MCFA intervenes in all the self-attention blocks of the SD UNet, by replacing the keys and values, that are originally computed from projections of the generating frame features, with the ones computed from the attended frames.
@@
We also propose Spatial-$\eta$ (line $12$), that is a novel technique that enables to choose, on a pixel-by-pixel basis, whether to use DDIM or DDPM as a sampling scheme. This enables the usage of DDPM in regions of the images where novel content should be created (for example, when a new part of an object is entering the scene), while using DDIM in the other regions to ensure consistency and determinism where the already-present content is just moving. Note that this spatial map $\eta^{f}$ can be obtained in multiple ways from the physical simulation. For example, $\eta^{f}$ can be set to $1$ in regions of the image where the flow is not well-defined (pointing outside of the image boundaries) or in regions where the optical flow field has discontinuities.
@@ 4.1 Experimental setting
In this section, we show examples of video generation based on different physics simulations: rigid body motion, fluid dynamics and multi-agent systems. Given an optical flow, we apply it on the SD latent space using MotionCraft (code is available at https://mezzelfo.github.io/MotionCraft/). Then, we compare our method to Text2Video-Zero [20] that, to the best of our knowledge, is the only diffusion-based zero-shot method for video generation.
@@ 4.5 Ablations
4.6 Additional Qualitative Results
Fig. 10 shows some additional results of MotionCraft. The first row shows a tree growing. This video was obtained using a simple constant outward-facing radial optical flow applied only on the foliage. Note that while the tree grows, its shadow evolution is coherent. As the input flow is zero in this part of the image, the shadow consistency is recovered only by Stable Diffusion. The last row show a video obtained by applying MotionCraft to SDXL. This shows the generalizability of MotionCraft to different diffusion models, with also different resolutions. Hence, MotionCraft is able to produce high-resolution videos with a high level of detail.
Figure 10: Additional videos from MotionCraft. The last row is obtained by applying MotionCraft to SDXL [23]
@@ 5 Conclusions
Acknowledgements
This publication is part of the project PNRR-NGEU which has received funding from the MUR - DM 352/2022. This work was partially supported by the European Union under the Italian National Recovery and Resilience Plan (NRRP) of NextGenerationEU, partnership on “Telecommunications of the Future” (PE00000001 - program “RESTART”).
@@ Appendix A Background
[…]
@@ Appendix B Extendend Ablation Study
In this section we show the remaining ablations for the scene Earth, and additional ablations on two new scenes: Dragons and Satellite Scan. The ablations for cross frame attention mechanism can be found in Figures 11 and 12. The ablations of the Spatial-$\eta$ are shown in Figures 13 and 14. Moreover, we also show the contribution of the inversion mechanism in Figures 15, 16, and 17.
@@ Appendix C Obstacles, Different Physics and Additional Visual Results
Figure 18: Fluid simulation: pouring drink. First row: Eulerian simulation performed with $\Phi$-flow; Second row: resulting optical flow; Third row: MotionCraft; Fourth row: resulting video when optical flow is applied directly to pixel-space.
Figure 19: Smoke simulation: Evaporating man. First and second row: Optical flow and video generated by MotionCraft with an Eulerian simulation. Third and fourth row: Optical flow and video generated by MotionCraft with a Lagrangian simulation.
@@ Appendix F Implementation Details and Licenses
We used the following hyperparameters throughout the work if not explicitely said otherwise. We set $\tau=400$, the number of inference steps (both for DDIM inversion and for inverse diffusion) is set to $200$ and the used model is runwayml/stable-diffusion-v1-5 (license CreativeML Open RAIL-M). All our experiments are done on a single NVIDIA A6000 (48GB); video generation runs in minutes (1-5min) on a single GPU. Our provided code is available under MIT license. The Earth image is a composite of six separate orbits taken on January 23, 2012 by the Suomi National Polar-orbiting Partnership satellite (Credit: NASA/NOAA).
@@ Appendix G Broader Impact
NeurIPS Paper Checklist
  1. 1.
    Claims
  2. Question: Do the main claims made in the abstract and introduction accurately reflect the paper’s contributions and scope?
  3. Answer: [Yes]
  4. Justification: All the claims in the abstract and introduction sections are supported by experimental evidence throughout the paper and reflect the results of our method.
  5. Guidelines:
    • The answer NA means that the abstract and introduction do not include the claims made in the paper.
    • The abstract and/or introduction should clearly state the claims made, including the contributions made in the paper and important assumptions and limitations. A No or NA answer to this question will not be perceived well by the reviewers.
    • The claims made should match theoretical and experimental results, and reflect how much the results can be expected to generalize to other settings.
    • It is fine to include aspirational goals as motivation as long as it is clear that these goals are not attained by the paper.
  6. 2.
    Limitations
  7. Question: Does the paper discuss the limitations of the work performed by the authors?
  8. Answer: [Yes]
  9. Justification: We discuss the limitations of the proposed work in the Appendix Section E.
  10. Guidelines:
    • The answer NA means that the paper has no limitation while the answer No means that the paper has limitations, but those are not discussed in the paper.
    • The authors are encouraged to create a separate "Limitations" section in their paper.
    • The paper should point out any strong assumptions and how robust the results are to violations of these assumptions (e.g., independence assumptions, noiseless settings, model well-specification, asymptotic approximations only holding locally). The authors should reflect on how these assumptions might be violated in practice and what the implications would be.
    • The authors should reflect on the scope of the claims made, e.g., if the approach was only tested on a few datasets or with a few runs. In general, empirical results often depend on implicit assumptions, which should be articulated.
    • The authors should reflect on the factors that influence the performance of the approach. For example, a facial recognition algorithm may perform poorly when image resolution is low or images are taken in low lighting. Or a speech-to-text system might not be used reliably to provide closed captions for online lectures because it fails to handle technical jargon.
    • The authors should discuss the computational efficiency of the proposed algorithms and how they scale with dataset size.
    • If applicable, the authors should discuss possible limitations of their approach to address problems of privacy and fairness.
    • While the authors might fear that complete honesty about limitations might be used by reviewers as grounds for rejection, a worse outcome might be that reviewers discover limitations that aren’t acknowledged in the paper. The authors should use their best judgment and recognize that individual actions in favor of transparency play an important role in developing norms that preserve the integrity of the community. Reviewers will be specifically instructed to not penalize honesty concerning limitations.
  11. 3.
    Theory Assumptions and Proofs
  12. Question: For each theoretical result, does the paper provide the full set of assumptions and a complete (and correct) proof?
  13. Answer: [N/A]
  14. Justification:
  15. Guidelines:
    • The answer NA means that the paper does not include theoretical results.
    • All the theorems, formulas, and proofs in the paper should be numbered and cross-referenced.
    • All assumptions should be clearly stated or referenced in the statement of any theorems.
    • The proofs can either appear in the main paper or the supplemental material, but if they appear in the supplemental material, the authors are encouraged to provide a short proof sketch to provide intuition.
    • Inversely, any informal proof provided in the core of the paper should be complemented by formal proofs provided in appendix or supplemental material.
    • Theorems and Lemmas that the proof relies upon should be properly referenced.
  16. 4.
    Experimental Result Reproducibility
  17. Question: Does the paper fully disclose all the information needed to reproduce the main experimental results of the paper to the extent that it affects the main claims and/or conclusions of the paper (regardless of whether the code and data are provided or not)?
  18. Answer: [Yes]
  19. Justification: We report a detailed pseudocode 1 from which it is possible to reproduce the main experimental results of our work. The method with all hyperparameters and implementation details are present in the Appendix F.
  20. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • If the paper includes experiments, a No answer to this question will not be perceived well by the reviewers: Making the paper reproducible is important, regardless of whether the code and data are provided or not.
    • If the contribution is a dataset and/or model, the authors should describe the steps taken to make their results reproducible or verifiable.
    • Depending on the contribution, reproducibility can be accomplished in various ways. For example, if the contribution is a novel architecture, describing the architecture fully might suffice, or if the contribution is a specific model and empirical evaluation, it may be necessary to either make it possible for others to replicate the model with the same dataset, or provide access to the model. In general. releasing code and data is often one good way to accomplish this, but reproducibility can also be provided via detailed instructions for how to replicate the results, access to a hosted model (e.g., in the case of a large language model), releasing of a model checkpoint, or other means that are appropriate to the research performed.
    • While NeurIPS does not require releasing code, the conference does require all submissions to provide some reasonable avenue for reproducibility, which may depend on the nature of the contribution. For example
      1. (a)
        If the contribution is primarily a new algorithm, the paper should make it clear how to reproduce that algorithm.
      2. (b)
        If the contribution is primarily a new model architecture, the paper should describe the architecture clearly and fully.
      3. (c)
        If the contribution is a new model (e.g., a large language model), then there should either be a way to access this model for reproducing the results or a way to reproduce the model (e.g., with an open-source dataset or instructions for how to construct the dataset).
      4. (d)
        We recognize that reproducibility may be tricky in some cases, in which case authors are welcome to describe the particular way they provide for reproducibility. In the case of closed-source models, it may be that access to the model is limited in some way (e.g., to registered users), but it should be possible for other researchers to have some path to reproducing or verifying the results.
  21. 5.
    Open access to data and code
  22. Question: Does the paper provide open access to the data and code, with sufficient instructions to faithfully reproduce the main experimental results, as described in supplemental material?
  23. Answer: [Yes]
  24. Justification: There is a link in the abstract pointing to the official project page, containing a public version of this code.
  25. Guidelines:
    • The answer NA means that paper does not include experiments requiring code.
    • Please see the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • While we encourage the release of code and data, we understand that this might not be possible, so “No” is an acceptable answer. Papers cannot be rejected simply for not including code, unless this is central to the contribution (e.g., for a new open-source benchmark).
    • The instructions should contain the exact command and environment needed to run to reproduce the results. See the NeurIPS code and data submission guidelines (https://nips.cc/public/guides/CodeSubmissionPolicy) for more details.
    • The authors should provide instructions on data access and preparation, including how to access the raw data, preprocessed data, intermediate data, and generated data, etc.
    • The authors should provide scripts to reproduce all experimental results for the new proposed method and baselines. If only a subset of experiments are reproducible, they should state which ones are omitted from the script and why.
    • At submission time, to preserve anonymity, the authors should release anonymized versions (if applicable).
    • Providing as much information as possible in supplemental material (appended to the paper) is recommended, but including URLs to data and code is permitted.
  26. 6.
    Experimental Setting/Details
  27. Question: Does the paper specify all the training and test details (e.g., data splits, hyperparameters, how they were chosen, type of optimizer, etc.) necessary to understand the results?
  28. Answer: [Yes]
  29. Justification: We specify the experimental setting in section 4.1 and in the Appendix F
  30. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The experimental setting should be presented in the core of the paper to a level of detail that is necessary to appreciate the results and make sense of them.
    • The full details can be provided either with the code, in Appendix, or as supplemental material.
  31. 7.
    Experiment Statistical Significance
  32. Question: Does the paper report error bars suitably and correctly defined or other appropriate information about the statistical significance of the experiments?
  33. Answer: [Yes]
  34. Justification: We report mean and std of the results in Table 1.
  35. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The authors should answer "Yes" if the results are accompanied by error bars, confidence intervals, or statistical significance tests, at least for the experiments that support the main claims of the paper.
    • The factors of variability that the error bars are capturing should be clearly stated (for example, train/test split, initialization, random drawing of some parameter, or overall run with given experimental conditions).
    • The method for calculating the error bars should be explained (closed form formula, call to a library function, bootstrap, etc.)
    • The assumptions made should be given (e.g., Normally distributed errors).
    • It should be clear whether the error bar is the standard deviation or the standard error of the mean.
    • It is OK to report 1-sigma error bars, but one should state it. The authors should preferably report a 2-sigma error bar than state that they have a 96% CI, if the hypothesis of Normality of errors is not verified.
    • For asymmetric distributions, the authors should be careful not to show in tables or figures symmetric error bars that would yield results that are out of range (e.g. negative error rates).
    • If error bars are reported in tables or plots, The authors should explain in the text how they were calculated and reference the corresponding figures or tables in the text.
  36. 8.
    Experiments Compute Resources
  37. Question: For each experiment, does the paper provide sufficient information on the computer resources (type of compute workers, memory, time of execution) needed to reproduce the experiments?
  38. Answer: [Yes]
  39. Justification: We specify the compute resource and time of execution in Appendix F
  40. Guidelines:
    • The answer NA means that the paper does not include experiments.
    • The paper should indicate the type of compute workers CPU or GPU, internal cluster, or cloud provider, including relevant memory and storage.
    • The paper should provide the amount of compute required for each of the individual experimental runs as well as estimate the total compute.
    • The paper should disclose whether the full research project required more compute than the experiments reported in the paper (e.g., preliminary or failed experiments that didn’t make it into the paper).
  41. 9.
    Code Of Ethics
  42. Question: Does the research conducted in the paper conform, in every respect, with the NeurIPS Code of Ethics https://neurips.cc/public/EthicsGuidelines?
  43. Answer: [Yes]
  44. Justification: The work respects the Code of Ethics.
  45. Guidelines:
    • The answer NA means that the authors have not reviewed the NeurIPS Code of Ethics.
    • If the authors answer No, they should explain the special circumstances that require a deviation from the Code of Ethics.
    • The authors should make sure to preserve anonymity (e.g., if there is a special consideration due to laws or regulations in their jurisdiction).
  46. 10.
    Broader Impacts
  47. Question: Does the paper discuss both potential positive societal impacts and negative societal impacts of the work performed?
  48. Answer: [Yes]
  49. Justification: We discuss the broader impact of the work in Appendix G
  50. Guidelines:
    • The answer NA means that there is no societal impact of the work performed.
    • If the authors answer NA or No, they should explain why their work has no societal impact or why the paper does not address societal impact.
    • Examples of negative societal impacts include potential malicious or unintended uses (e.g., disinformation, generating fake profiles, surveillance), fairness considerations (e.g., deployment of technologies that could make decisions that unfairly impact specific groups), privacy considerations, and security considerations.
    • The conference expects that many papers will be foundational research and not tied to particular applications, let alone deployments. However, if there is a direct path to any negative applications, the authors should point it out. For example, it is legitimate to point out that an improvement in the quality of generative models could be used to generate deepfakes for disinformation. On the other hand, it is not needed to point out that a generic algorithm for optimizing neural networks could enable people to train models that generate Deepfakes faster.
    • The authors should consider possible harms that could arise when the technology is being used as intended and functioning correctly, harms that could arise when the technology is being used as intended but gives incorrect results, and harms following from (intentional or unintentional) misuse of the technology.
    • If there are negative societal impacts, the authors could also discuss possible mitigation strategies (e.g., gated release of models, providing defenses in addition to attacks, mechanisms for monitoring misuse, mechanisms to monitor how a system learns from feedback over time, improving the efficiency and accessibility of ML).
  51. 11.
    Safeguards
  52. Question: Does the paper describe safeguards that have been put in place for responsible release of data or models that have a high risk for misuse (e.g., pretrained language models, image generators, or scraped datasets)?
  53. Answer: [N/A]
  54. Justification:
  55. Guidelines:
    • The answer NA means that the paper poses no such risks.
    • Released models that have a high risk for misuse or dual-use should be released with necessary safeguards to allow for controlled use of the model, for example by requiring that users adhere to usage guidelines or restrictions to access the model or implementing safety filters.
    • Datasets that have been scraped from the Internet could pose safety risks. The authors should describe how they avoided releasing unsafe images.
    • We recognize that providing effective safeguards is challenging, and many papers do not require this, but we encourage authors to take this into account and make a best faith effort.
  56. 12.
    Licenses for existing assets
  57. Question: Are the creators or original owners of assets (e.g., code, data, models), used in the paper, properly credited and are the license and terms of use explicitly mentioned and properly respected?
  58. Answer: [Yes]
  59. Justification: We cite all the used assets throughout the work and we report the correct version used for our experiments.
  60. Guidelines:
    • The answer NA means that the paper does not use existing assets.
    • The authors should cite the original paper that produced the code package or dataset.
    • The authors should state which version of the asset is used and, if possible, include a URL.
    • The name of the license (e.g., CC-BY 4.0) should be included for each asset.
    • For scraped data from a particular source (e.g., website), the copyright and terms of service of that source should be provided.
    • If assets are released, the license, copyright information, and terms of use in the package should be provided. For popular datasets, paperswithcode.com/datasets has curated licenses for some datasets. Their licensing guide can help determine the license of a dataset.
    • For existing datasets that are re-packaged, both the original license and the license of the derived asset (if it has changed) should be provided.
    • If this information is not available online, the authors are encouraged to reach out to the asset’s creators.
  61. 13.
    New Assets
  62. Question: Are new assets introduced in the paper well documented and is the documentation provided alongside the assets?
  63. Answer: [Yes]
  64. Justification: We release the code along with the documentation needed to run the experiments.
  65. Guidelines:
    • The answer NA means that the paper does not release new assets.
    • Researchers should communicate the details of the dataset/code/model as part of their submissions via structured templates. This includes details about training, license, limitations, etc.
    • The paper should discuss whether and how consent was obtained from people whose asset is used.
    • At submission time, remember to anonymize your assets (if applicable). You can either create an anonymized URL or include an anonymized zip file.
  66. 14.
    Crowdsourcing and Research with Human Subjects
  67. Question: For crowdsourcing experiments and research with human subjects, does the paper include the full text of instructions given to participants and screenshots, if applicable, as well as details about compensation (if any)?
  68. Answer: [N/A]
  69. Justification:
  70. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Including this information in the supplemental material is fine, but if the main contribution of the paper involves human subjects, then as much detail as possible should be included in the main paper.
    • According to the NeurIPS Code of Ethics, workers involved in data collection, curation, or other labor should be paid at least the minimum wage in the country of the data collector.
  71. 15.
    Institutional Review Board (IRB) Approvals or Equivalent for Research with Human Subjects
  72. Question: Does the paper describe potential risks incurred by study participants, whether such risks were disclosed to the subjects, and whether Institutional Review Board (IRB) approvals (or an equivalent approval/review based on the requirements of your country or institution) were obtained?
  73. Answer: [N/A]
  74. Justification:
  75. Guidelines:
    • The answer NA means that the paper does not involve crowdsourcing nor research with human subjects.
    • Depending on the country in which research is conducted, IRB approval (or equivalent) may be required for any human subjects research. If you obtained IRB approval, you should clearly state this in the paper.
    • We recognize that the procedures for this may vary significantly between institutions and locations, and we expect authors to adhere to the NeurIPS Code of Ethics and the guidelines for their institution.
    • For initial submissions, do not include any information that would break anonymity (if applicable), such as the institution conducting the review.